hallm:

At 25°C, the screech of the Sneetches can travel 250 meters across the plains of Sneedville in 0.50 seconds. Determine the velocity of these sound waves as they travel across the plains.

Shadow:

\[V = \frac{ \Delta x }{ \Delta t }\] Velocity is equal to the change in displacement over the change in time. A dog traveled 10m in 2 seconds. \[V_{dog} = \frac{ 10 }{ 2 } = 5 m/s\]

hallm:

ohhhhhhh ok

hallm:

wait ok so you divide 10 divided by two and you get 5 but what next

Shadow:

That is it. Divide the displacement by time, then give the correct units, then you're done.

Shadow:

10/5 = 2 is just my example though. See how I did it then look at how to do yours.